OpenStudy (anonymous):

evlature the expression 7c6/7p6.

OpenStudy (anonymous):

whats the difference between p and c?

OpenStudy (anonymous):

7!/6! * (1/7!) = 1/6! =1/720

OpenStudy (anonymous):

and that's to lowest terms?

OpenStudy (anonymous):

a c b = a!/(b!*(a-b)! a p b = a!/(a-b)!

OpenStudy (anonymous):

would 1/720 be the lowest you can put it?

OpenStudy (anonymous):

Yes